In an effusion experiment, it was determined that nitrogen gas, N2, effused at a rate 1.812 times faster than an unknown gas. What is the molar mass of the unknown gas?
Answer #1

r1/r2=(M2/M1)^1/2

r1=r

r2 of N2=1.812r

M2=28

M1=?

r/1.812r=(28/M1)^1/2

1/1.812=(28/M2)^1/2

M2= 92.10 gram/mole
Answer #2

rate of effusion of a gas is inversely proportional to the square root of the molar mass of the gas

therefore, let the molar mass of unknown gas be M g/mole

now, rate = flow, per unit time

now, (rate of N2)/(rate of unknown gas) = sqrt(molar mass of unknown gas/ molar mass of N2)

or, 1.812 = sqrt(M/28)

or, M = 28*(1.812)^2 = 91.934 g
Answer #3

according to graham law

r=rate of effusio
rGas A/rGas B = (MGas B)